### v2

@@ 5 Results
In this section, we present the results of the experiments conducted to evaluate the proposed approach. In the first section, we compare the layers of the coordination network, while in the second part we analyze the different ways of operationalizing the multimodal coordinated behavior.
@@ 5.1.1 Layers comparison
Table 2: Statistics of the coordination network after applying the filtering steps.
weight statistics
@@
approach mean median stdDev max min nNodes nEdges nCC nCom
monomodal RTW 0.189 0.171 0.060 1 0.126 $17,891$ $1,760,094$ 64 100
RPL 0.610 0.584 0.121 1.000 0.451 $16,071$ $1,334,039$ 6 23
URL 0.923 0.936 0.074 1.000 0.791 $18,741$ $2,302,080$ 117 134
MEN 0.173 0.151 0.072 1.000 0.111 $18,444$ $3,047,793$ 61 95
HST 0.362 0.335 0.101 1.000 0.246 $18,300$ $1,520,067$ 126 186
multimodal UNFL (nw) - - - - - $39,646$ $7,303,637$ 106 175
UNFL (ec) 1.364 1.000 0.688 5.000 1.000 $39,646$ $7,303,637$ 106 154
UNFL (sum) 0.596 0.537 0.371 3.971 0.111 $39,646$ $7,303,637$ 106 169
MULTI - - - - - - - - 180
INTFL 2.245 2.219 0.235 3.849 1.826 257 375 19 27
@@
$\dagger$ nCC: number of connected components, $\ddagger$ nCom: number of communities
@@
Statistics comparison. In Table 2, we report key statistics for each implemented method, including edge weight statistics, the number of nodes, edges, connected components, and communities. UNFL (nw) contains empty values for the weight statistics because the flattened network is unweighted. Similarly, the MULTI approach also presents mostly empty values across the statistics, as it operates on a multiplex network composed of five separate layers – meaning that the relevant statistics are computed individually for each layer. Among the layers, URL and RPL exhibit the highest mean edge weights, while MEN has the lowest. RTW has the highest number of edges, whereas MEN has the fewest. Interestingly, RPL has only six connected components, with one large connected component and a few small isolated ones. However, for all layers, a large connected component exists, which is the primary focus of our analysis. In terms of community count, the layers show similar results, except for HST, which has the highest number of communities, and RPL, which has the lowest.
@@
Therefore, the upper part of Table 2 compares different single-layer approaches, while the lower part presents the results of the multimodal approaches. This analysis offers an initial comparison of the different operationalizations of multimodality. Among them, INTFL results in the lowest number of nodes and edges, whereas UNFL captures the highest. INTFL significantly reduces the network size since it retains only edges present in all layers, identifying the fewest communities. In addition, the mean edge weight is significantly higher with respect to the other flattening approaches. On one hand, this highlights that INTFL captures extremely significant coordinated behaviors, but on the other hand, it is very selective and it is the strictest operationalization of multimodality, limiting the amount of captured information compared to other methods. For this reason, our focus will shift to comparing INDI, UNFL, and MULTI approaches.
(a) Actor coverage.
(b) Edge coverage.
(c) Pearson degree correlation.
@@
Figure 4: Metrics comparison between the layers of the coordination network. The computed metrics are on the actors and edges coverage, Jaccard similarity, Jeffrey divergence, and Pearson correlation.
@@
Metrics comparison. In this analysis, we leverage metrics recommended by Bródka et al. [68] for evaluating the layers of a multiplex network: actor coverage, edge coverage, and Pearson correlation. Figure 4 presents heatmaps comparing the layers, or modalities, of the coordination network. Each heatmap displays a pairwise comparison between two layers, with rows and columns representing the layers and cells containing the computed metrics.
@@
Actor coverage. The coverage metrics assess the percentage of actors and edges shared between two layers, offering insights into the degree of overlap. These metrics are directional, meaning that the coverage $cov(i,j)\neq cov(j,i)$. Specifically, the value in cell $(i,j)$ represents the percentage of actors or edges in layer $i$ that are also present in layer $j$. The coverage values range from 00 to $1$, where $1$ indicates complete overlap between the actors or edges in layer $i$ and those in layer $j$.
@@
Results, showed in Figure 4(a), indicates that MEN has the largest values with both RTW and HST in both directions. This implies that the actors involved in MEN are largely present in RTW and HST, and vice versa. Conversely, RPL exhibits the smallest actor coverage with all other layers, signifying that actors in RPL are less involved in the other layers. URL, excluding RPL, shows a moderate level of actor coverage with all other layers. These findings suggest that RPL is the most distinct layer in terms of actor involvement, while RTW and MEN appear to be quite similar. In particular, the actor coverage values are comparable in both directions, reflecting the similar number of actors present in these two layers.
@@
Edge coverage. Results, showed in Figure 4(b), mirror those of actor coverage but exhibit some directional variations. For example, RTW has an edge coverage of $0.707$ with MEN, while MEN has only $0.409$ with RTW. This disparity arises because co-mention contains twice as many edges as RTW, leading to greater representation of RTW edges within MEN. HST also shares a significant proportion of edges with MEN. In contrast, RPL and URL show the smallest edge coverage with all other layers. For RPL, this is expected due to its low actor coverage. However, the small edge coverage of URL is surprising, given its moderate actor coverage. This discrepancy might result from the unique nature of URL-sharing actions compared to other actions.
@@
Pearson degree correlation. The Pearson correlation evaluates the linear relationship between the degree of actors in two layers, with values ranging from $-1$ to $1$. A correlation of $-1$ indicates that high-degree actors in one layer are low-degree in the other and vice versa, while a correlation of $1$ indicates that high-degree (or low-degree) actors in one layer are also high-degree (or low-degree) actors in the other. Notably, this metric focuses on the number of incident edges for each actor-layer pair, without considering the specific adjacent actors.
@@
Results, showed in Figure 4(c), reveal a strong similarity in actor degree distribution among RTW, MEN, and HST, with the strongest correlation observed between RTW and MEN. In contrast, the weakest correlation is between RPL and URL, further highlighting their distinctiveness.
@@
These preliminary findings highlight that RPL, and to a lesser extent URL, stand out as distinct layers, while RTW, MEN, and HST demonstrate considerable similarity. This result underscores the complementary nature of the different layers: although they provide diverse information, actors tend to maintain similar roles across layers.
@@ 5.1.2 Communities comparison
Normalized Mutual Information. We begin with a quantitative analysis to evaluate the similarity between the communities identified by different monomodal approaches. To achieve this, we use Normalized Mutual Information ($NMI$), a metric designed to compare two partitions of a set of nodes [69, 70, 71]. $NMI$ quantifies the extent to which the community structures from different modalities align, providing a standardized measure of agreement between the partitions. This analysis allows us to assess how closely the community detection results from different monomodal approaches correspond to other ones. The $NMI$ is computed considering the communities with more than 200 nodes for all couples of co-actions, and the results are reported in Figure 5. This analysis confirms that RTW and MEN are quite similar in terms of communities detected, reaching the maximum value equal to $NMI=0.38$. The smallest values are reached by all co-actions with RPL, which is in line with the previous results, showing that RPL is the most different co-action in terms of detected communities.
@@
Figure 5: Normalized Mutual Information ($NMI$) heatmap between the set of communities with size greater than 200 for each couple of co-actions.
(a) RTW vs. INDI Lost, common, and gained communities.
(b) RTW vs. INDI Lost, common, and gained nodes.
@@
Figure 6: Bar charts showing the number of the lost, common, and gained communities and nodes switching from INDI to RTW.
@@
Lost, common, and gained communities and nodes. Based on the methodology described in Section 3.3.1, we compute the overlapping of all possible combinations of communities, with size greater than 200 nodes. The matrices obtained from the analysis can be effectively visualized as heatmaps, providing an intuitive representation of the underlying relationships. Appendix A reports the heatmaps for the performed comparisons. Specifically, to answer the first research question RTW vs. INDI, and the second research question UNFL vs. INDI and MULTI vs. INDI. A comparison of the overlapping matrices of the three UNFL strategies reveals minimal differences among them. Therefore, we will focus on the UNFL (sum) strategy in subsequent analyses, as it retains the most information. Given the large number of results to present, we report only the aggregated outcomes for brevity.
@@
Hence, following the methodology defined in Section 3.3.2, we analyze the overlapping matrices of the communities identified by the different approaches, by labeling the communities as lost, common, and gained switching from one approach to another one. We set the threshold $\theta=0.5$, but we obtain similar results for $\theta=0.7$ and $\theta=0.9$. A higher threshold reduces the number of common communities, but the patterns observed remain still valid. We focus only on RTW vs. INDI, since it is just a matter of perspective, and RTW is the most common and studied action in the literature [3, 18, 26, 72, 73]. Similarly, we analyze the lost, common, and gained nodes, according to the methodology defined in Section 3.3.3. Figure 6 presents bar charts summarizing how the communities and nodes identified in the INDI approaches correspond to those found in the RTW network. Specifically, communities and nodes are classified as lost, common, or gained, based on whether they are no longer detected, consistently identified, or newly discovered in RTW compared to the INDI approaches. These results allows to compare the different MONO approaches, allowing to understand if different co-actions provide different information.
@@
Figure 6(a) highlights that RPL and RTW capture completely different structures: they share no communities in common, with RTW identifying five new communities and losing the two detected by RPL. Similarly, URL and RTW exhibit almost entirely distinct results, as RTW loses seven communities compared to URL and identifies four new ones. In contrast, RTW and MEN fully overlap, detecting the same set of coordinated communities and thus conveying identical information about user coordination. Finally, RTW and HST share four communities, while three communities are detected exclusively by HST. This indicates that RTW and HST partially capture complementary aspects of the coordinated behavior, while still identifying some common structures.
@@
Similarly, Figure 6(b) shows that RTW differs substantially from both RPL and URL in terms of the coordinated nodes it detects. However, when shifting the focus from communities to individual nodes, no MONO approach clearly emerges as more informative. Rather, the information gained by RTW is approximately balanced by what it loses, indicating that each modality captures distinct yet equally important facets of coordination. Consistent with previous findings, RTW remains strongly aligned with MEN, while a similar but weaker alignment is observed with HST.
@@ 5.1.3 Common communities characterization
While we have compared the community results across modalities in terms of node overlap, this does not ensure that the position or role of these communities within the network remains consistent across layers. Even if the same users are identified as part of the communities, their structural properties and interactions may vary depending on the modality. To address this, we compute a set of metrics to characterize the communities previously labeled as common across modalities. This analysis allows us to assess whether these communities exhibit similar structural and functional characteristics across different layers, providing deeper insights into the interplay between modalities. To conduct this analysis, we compute a set of metrics for each community identified as common across modalities (i.e., detected by both RTW and another MONO approach). The metrics include size, density, average degree, average weight, average clustering coefficient, conductance, and assortativity [74]. These metrics provide a comprehensive characterization of the communities, capturing their structural and functional properties within the network. This approach allows us to evaluate and compare the communities’ characteristics across different modalities systematically.
@@
Table 3: Cosine similarity between the vectors of the communities in common with RTW communities. The last column reports the average cosine similarity for each layer.
$layer_{1}$ $layer_{2}$
layer community layer community cosine similarity mean
RTW 0 MEN 0 0.993 0.971
RTW 1 MEN 1 0.952
RTW 2 MEN 2 0.969
RTW 3 MEN 3 0.968
RTW 4 MEN 4 0.976
RTW 0 HST 0 0.961 0.923
RTW 1 HST 2 0.836
RTW 2 HST 3 0.982
RTW 4 HST 5 0.915
RTW 1 URL 4 0.537 0.537
@@
For each pair of common labeled communities (as identified in Figure 6(a)), we calculate the cosine similarity between their metrics. Higher cosine similarity values indicate greater similarity between the communities in terms of their structural and functional properties. Additionally, we compute the mean cosine similarity for the common communities within each layer. The results, presented in Table 3, reveal that MEN and HST layers exhibit very high levels of cosine similarity, suggesting strong alignment in the structural and functional characteristics of the detected communities. This aligns with the finding that RTW shares a substantial number of communities with both MEN and HST, reinforcing the idea that these modalities capture closely related coordination patterns. In contrast, the URL layer not only has very few communities in common with RTW, but the only shared community also exhibits significant structural differences. This suggests that URL captures a distinct form of coordination, diverging more from retweet-based interactions compared to mention and hashtag.
@@
A more detailed representation of the common communities similarity is presented by Figure 7, which visualizes as a starplot the normalized scores between 0 and 1 of the aforementioned metrics for three examples of couple of common community. The starplot are color-coded according to the modality to which each community belongs, with a distinct color representing each MONO approach. This visualization allows us to compare the metric scores of matched communities and assess the degree of similarity between them. If two modalities convey similar information, the matched communities will exhibit comparable metric scores, resulting in similar starplot shapes. Conversely, differences in the shapes of the radar plots indicate variations in the information captured by the modalities. This representation effectively highlights the similarities and discrepancies across modes in a clear and intuitive manner. As Figure 6(a) shows, RPL and RTW do not have communities in common, and for this reason there is no starplot for comparison. MEN common communities exhibit similar values across all metrics, whereas HST communities are characterized by a high average weight, and their starplot shape slightly differs from that of RTW. Lastly, URL, which shares only one common community with RTW, presents a completely distinct starplot shape, with notably high values for both average weight and assortativity.
(a) RTW vs. MEN
(b) RTW vs. HST
(c) RTW vs. URL
@@
Figure 7: Example of metrics comparison of three couple of common communities between RTW and MEN, HST, and URL.
@@
Figure 8: Visualization of common communities between RTW and other co-actions in a two-dimensional space using PCA. Each community is represented by a vector of metrics, i.e., size, density, average degree, average weight, average clustering coefficient, conductance, and assortativity. These high dimensional vectors are projected into a two-dimensional space using PCA. Communities are represented with distinct point shapes based on the co-action and are linked to their corresponding RTW community. Each RTW community may be associated with multiple communities from different co-actions, with all connected communities sharing the same color for clarity.
@@
The starplot visualization allows to have a quick understanding of the differences between the communities, but it is very detailed and does not provide a global view of the differences. Hence, we represent each community as a feature vector of the aforementioned metrics. Then the feature vectors are projected into a two-dimensional space using Principal Component Analysis (PCA). Figure 8 shows the scatterplot in two PCA dimensions, where each community is linked with the one in common with RTW and identified with a different point shape according to the co-action. Each RTW community may be associated with multiple communities from different co-actions, and the cluster of connected communities are depicted with the same color. The PCA visualization highlights how each cluster of common communities – same color – occupies a distinct region of the plot. This spatial separation suggests that common communities of different INDI are structurally similar to each other while remaining distinct from non-common communities. This reinforces the idea that coordination patterns within common communities are preserved across different modalities. However, an exception emerges with the URL layer, which appears more dispersed and does not conform to the same clustering pattern. This aligns with previous findings that URL is the most distinct and unconventional layer, capturing a different form of coordination compared to the others.

### v3

@@ 5 Results
In this section, we present the results of the experiments conducted to evaluate the proposed approach. In the first section, we compare the layers of the coordination network, while in the second part, we analyze the different ways of operationalizing the multimodal coordinated behavior.
@@ 5.1.1 Layers comparison
Table 3: Statistics of the UK coordination network after applying the filtering steps.
network’s weight statistics
@@
approach mean median stdDev max min nNodes nEdges nCC nCLV nCIM‡‡
monomodal RTW 0.189 0.171 0.060 1.000 0.126 $17,891$ $1,760,094$ 64 100 65
RPL 0.610 0.584 0.121 1.000 0.451 $16,071$ $1,334,039$ 6 23 211
URL 0.923 0.936 0.074 1.000 0.791 $18,741$ $2,302,080$ 117 134 351
MEN 0.173 0.151 0.072 1.000 0.111 $18,444$ $3,047,793$ 61 95 145
HST 0.362 0.335 0.101 1.000 0.246 $18,300$ $1,520,067$ 126 186 127
multimodal UNFL (nw) - - - - - $39,646$ $7,303,637$ 106 175 366
UNFL (ec) 1.364 1.000 0.688 5.000 1.000 $39,646$ $7,303,637$ 106 154 382
UNFL (sum) 0.596 0.537 0.371 3.971 0.111 $39,646$ $7,303,637$ 106 169 425
MULTI - - - - - - - - 180 124
INTFL 2.245 2.219 0.235 3.849 1.826 257 375 19 27 43
@@
$\dagger$ nCC: number of connected components, $\ddagger$ nComLV: number of communities with Louvain
$\ddagger\ddagger$ nComIM: number of communities with Infomap
Table 4: Statistics of the IORussia coordination network after applying the filtering steps.
network’s weight statistics
approach mean median stdDev max min nNodes nEdges nCC nCLV nCIM‡‡
monomodal RTW 0.326 0.236 0.233 1.000 0.137 $3,397$ $27,566$ 441 453 441
RPL 0.774 0.750 0.132 1.000 0.584 $203$ $826$ 18 21 26
URL 0.534 0.495 0.208 1.000 0.268 $2,327$ $5,698$ 450 455 449
MEN 0.251 0.191 0.180 1.000 0.117 $4,874$ $49,478$ 663 679 663
HST 0.431 0.362 0.187 1.000 0.246 $3,169$ $33,512$ 348 361 348
multimodal UNFL (nw) - - - - - $6,921$ $77,972$ 762 785 761
UNFL (ec) 1.502 1.000 0.721 5.000 1.000 $6,921$ $77,972$ 762 785 762
UNFL (sum) 0.507 0.408 0.410 4.000 0.117 $6,921$ $77,972$ 762 790 763
MULTI - - - - - - - - 798 762
INTFL 1.621 1.617 0.058 1.680 1.564 6 3 19 3 3
$\dagger$ nCC: number of connected components, $\ddagger$ nComLV: number of communities with Louvain
$\ddagger\ddagger$ nComIM: number of communities with Infomap
@@
Statistics comparison. Table 3 and Table 4 respectively for UK and IORussia report key statistics for each implemented method, including edge weight statistics, the number of nodes, edges, connected components, and communities. UNFL (nw) contains empty values for the weight statistics because the flattened network is unweighted. Similarly, the MULTI approach also presents mostly empty values across the statistics, as it operates on a multiplex network composed of five separate layers – meaning that the relevant statistics are computed individually for each layer. In both datasets, among the layers, URL and RPL exhibit the highest mean edge weights, while MEN has the lowest. MEN has the highest number of edges, whereas RPL has the fewest. Interestingly, RPL has only 6 (UK) and 18 (IORussia) connected components, with one large connected component and a few small isolated ones. However, for all layers, a large connected component exists.
The similar number of communities detected by Louvain and Infomap on IORussia dataset is likely due to its more strongly coordinated structure: state-linked IO campaigns tend to form dense, well-separated clusters, which may lead different detection strategies to converge on comparable partitions. In contrast, the UK dataset is more heterogeneous, with diverse user types and short-lived interaction bursts during the election period. This weaker structural cohesion likely causes Louvain to merge loosely connected regions, while Infomap captures finer-grained distinctions and consequently identifies many more communities, though we cannot be certain that this is the only factor driving the difference.
@@
Therefore, in both datasets, the upper part of the tables compares different single-layer approaches, while the lower part presents the results of the multimodal approaches. This analysis offers an initial comparison of the different operationalizations of multimodality. Among them, INTFL results in the lowest number of nodes and edges, whereas UNFL captures the highest. INTFL significantly reduces the network size since it retains only edges present in all layers, identifying the fewest communities. In addition, the mean edge weight is significantly higher with respect to the other flattening approaches. On one hand, this suggests that INTFL may capture extremely clear coordinated behaviors, exploiting all modalities at the same time. On the other hand, it is very selective, and it is the strictest operationalization of multimodality, limiting the amount of captured information compared to other methods. For this reason, our focus will shift to comparing INDE, UNFL, and MULTI approaches.
@@
Figure 4: Metrics comparison between the layers of the UK and Russia coordination network. The computed metrics are on the actors and edges coverage, and Pearson correlation.
@@
Metrics comparison. In this analysis, we leverage metrics recommended by Bródka et al. [48] for comparing the layers of a multiplex network: actor coverage, edge coverage, and Pearson correlation. Figure 4 presents heatmaps comparing the layers, or modalities, of the coordination network. Each heatmap displays a pairwise comparison between two layers, with rows and columns representing the layers and cells containing the computed metrics.
@@
Actor coverage. The coverage metrics assess the percentage of actors and edges shared between two layers, offering insights into the degree of overlap. These metrics are directional, meaning that the coverage $cov(i,j)\neq cov(j,i)$. Specifically, the value in cell $(i,j)$ represents the percentage of actors or edges in layer $i$ that are also present in layer $j$. The coverage values range from 0 to $1$, where $1$ indicates complete overlap between the actors or edges in layer $i$ and those in layer $j$.
@@
Rnoesults for UK, shown in Figure 4a, indicate that MEN has the largest values with both RTW and HST in both directions. This implies that the actors involved in MEN are largely present in RTW and HST, and vice versa. Conversely, RPL exhibits the smallest actor coverage with all other layers, signifying that actors in RPL are less involved in the other layers. URL, excluding RPL, shows a moderate level of actor coverage with all other layers. These findings suggest that RPL is the most distinct layer in terms of actor involvement, while RTW and MEN appear to be quite similar. In particular, the actor coverage values are comparable in both directions, reflecting the similar number of actors present in these two layers.
Instead, results for IORussia, depicted in Figure 4d, show that RPL has high values of coverage with all the layers, except URL, meaning that most of the actors involved in the other layers are also present in RPL. But this was expected, since RPL is very small in terms of the number of actors (see Table 4). Indeed, in the other direction, RPL has very small values of coverage with all the layers. Also MEN have high values of coverage in one direction with all the other layers. As for UK, MEN and RTW still exhibit high actor coverage in both directions, indicating a strong overlap of actors between these two layers.
@@
Edge coverage. Results for UK, shown in Figure 4b, mirror those of actor coverage but exhibit some directional variations. For example, RTW has an edge coverage of $0.707$ with MEN, while MEN has only $0.409$ with RTW. This disparity arises because co-mention contains twice as many edges as RTW, leading to greater representation of RTW edges within MEN. HST also shares a significant proportion of edges with MEN. In contrast, RPL and URL show the smallest edge coverage with all other layers. For RPL, this is expected due to its low actor coverage. However, the small edge coverage of URL is surprising, given its moderate actor coverage. This discrepancy might result from the unique nature of URL-sharing actions compared to other actions. Results for IORussia, shown in Figure 4e, are quite similar to those of UK. Indeed, RTW and RPL have high values of edge coverage with MEN, but not vice versa, due to the different number of edges in the layers.
@@
Pearson degree correlation. The Pearson correlation evaluates the linear relationship between the degree of actors in two layers, with values ranging from 1-1 to $1$. A correlation of 1-1 indicates that high-degree actors in one layer are low-degree in the other and vice versa, while a correlation of $1$ indicates that high-degree (or low-degree) actors in one layer are also high-degree (or low-degree) actors in the other. Notably, this metric focuses on the number of incident edges for each actor-layer pair, without considering the specific adjacent actors.
@@
UK results, shown in Figure 4c, reveal a strong similarity in actor degree distribution among RTW, MEN, and HST, with the strongest correlation observed between RTW and MEN. In contrast, the weakest correlation is between RPL and URL, further highlighting their distinctiveness. Similarly, IORussia results, shown in Figure 4f, indicate that RTW, MEN, and HST maintain a strong correlation in actor degree distribution, with RTW and MEN again exhibiting the highest correlation.
@@
These preliminary findings highlight on both datasets that RPL, and to a lesser extent URL, stand out as distinct layers, while RTW, MEN, and HST show considerable similarity. This result underscores the complementary nature of the different layers: although they provide diverse information, actors tend to maintain similar roles across layers.
@@ 5.1.2 Communities comparison
Figure 5: Normalized Mutual Information (NMI) heatmap comparing the communities detected with the Louvain algorithm, considering only those with size greater than 200 for UK, and 70 for IORussia, across all pairs of co-actions.
Normalized Mutual Information. We begin with a quantitative analysis to evaluate the similarity between the communities identified by different monomodal approaches. To achieve this, we use Normalized Mutual Information ($NMI$), a metric designed to compare two partitions of a set of nodes [49]. $NMI$ quantifies the extent to which the community structures from different modalities align, providing a standardized measure of agreement between the partitions. This analysis allows us to assess how closely the community detection results from different monomodal approaches correspond to each other. From now on, we consider the communities with more than 200 nodes for UK and 70 nodes for IORussia, for all couples of co-actions, and results obtained with Infomap are reported in Appendix A for completeness.
Results for both datasets are reported in Figure 5. Specifically, Figure 5a shows results for UK, which confirm that RTW and MEN are quite similar in terms of communities detected, reaching the maximum value equal to $NMI=0.38$. The smallest values are reached by all co-actions with RPL, which is in line with the previous results, showing that RPL is the most different co-action in terms of detected communities. IORussia results, shown in Figure 5b, present high values of NMI between all combinations of co-actions, except those between RPL and URL, which are lower. In general, the NMI values for IORussia are higher than those for UK, indicating a greater similarity in community structures across modalities in IORussia. This suggests that coordinated behavior in IORussia is more consistently captured across different modalities compared to the UK.
@@
Figure 6: Bar charts for Louvain approach showing the number of lost, common, and gained communities and nodes switching from INDE to RTW.
@@
Lost, common, and gained communities and nodes. Based on the methodology described in Section 3.3.1, we compute the overlapping of all possible combinations of communities, with size greater than 200 nodes for UK and 70 nodes for IORussia. The matrices obtained from the analysis can be effectively visualized as heatmaps, providing an intuitive representation of the underlying relationships. Appendix B and Appendix C present, for UK, the heatmaps corresponding to the comparisons performed on the communities detected using the Louvain and Infomap approaches, respectively. Similarly, Appendix D and Appendix E report the corresponding results for IORussia. Specifically, to answer the first research question RTW vs. INDE, and the second research question UNFL vs. INDE and MULTI vs. INDE. A comparison of the overlapping matrices of the three UNFL strategies reveals minimal differences among them. Therefore, we will focus on the UNFL (sum) strategy in subsequent analyses, as it retains the most information. Given the large number of results to present, we report only the aggregated outcomes for brevity.
@@
Hence, following the methodology defined in Section 3.3.2, we analyze the overlapping matrices of the communities identified by the different approaches, by labeling the communities as lost, common, and gained switching from one approach to another one. We set the threshold $\theta=0.5$, but we obtain similar results for $\theta=0.7$ and $\theta=0.9$. A higher threshold reduces the number of common communities, but the patterns observed remain still valid. We focus only on RTW vs. INDE, as labeling communities with respect to different co-actions is simply a change in perspective, and RTW is the most common and widely studied action in the literature [15]. Similarly, we analyze the lost, common, and gained nodes, according to the methodology defined in Section 3.3.3. Figure 6 presents bar charts summarizing how the communities and nodes identified in the INDE approaches correspond to those found in the RTW network. Specifically, communities and nodes are classified as lost, common, or gained, based on whether they are no longer detected, consistently identified, or newly discovered in RTW compared to the INDE approaches. These results compare the different MONO approaches, allowing to understand if different co-actions provide different information.
@@
Results of UK with Louvain approach are shown in Figure 6a, which highlights that RPL and RTW capture completely different structures: they have no communities in common, with RTW identifying five new communities and losing the two detected by RPL. Similarly, URL and RTW exhibit almost entirely distinct results, as RTW loses seven communities compared to URL and identifies four new ones. In contrast, RTW and MEN fully overlap, detecting the same set of coordinated communities and thus conveying identical information about user coordination. Finally, RTW and HST share four communities, while three communities are detected exclusively by HST. This indicates that RTW and HST partially capture complementary aspects of the coordinated behavior, while still identifying some common structures.
@@
Similarly, Figure 6b shows that RTW differs substantially from both RPL and URL in terms of the coordinated nodes it detects. However, when shifting the focus from communities to individual nodes, no MONO approach clearly emerges as more informative. Rather, the information gained by RTW is approximately balanced by what it loses, indicating that each modality captures distinct yet equally important facets of coordination. Consistent with previous findings, RTW remains strongly aligned with MEN, while a similar but weaker alignment is observed with HST.
IORussia results for communities and nodes are shown in Figure 6c and 6d, respectively. Both with Louvain and Infomap approach, RTW shares some communities in common with MEN and HST, while it does not share any community with RPL. Finally, Appendix A reports results for Infomap approach. Figure A.2 shows similar results for Infomap approach on UK, with the only difference that RTW and MEN do not share all the communities in common, but RTW loses three communities with respect to MEN.
Both datasets and both community detection approaches consistently highlight that RPL is a co-action that differs substantially from RTW, whereas MEN and HST tend to share more information. However, no strong or systematic overlap emerges across the different co-actions. This implies that none of them can be excluded a priori from a multimodal analysis: each co-action may contribute complementary information about coordinated behavior.
@@ 5.1.3 Common communities characterization
While we have compared the community results across modalities in terms of node overlap, this does not ensure that the position or role of these communities within the network remains consistent across layers. Even if the same users are identified as part of the communities, their structural properties and interactions may vary depending on the modality. To address this, we compute a set of metrics to characterize the communities previously labeled as common across modalities. This analysis allows us to assess whether these communities exhibit similar structural and functional characteristics across different layers, providing deeper insights into the interplay between modalities. To conduct this analysis, we compute a set of metrics for each community identified as common across modalities (i.e., detected by both RTW and another MONO approach). The metrics include size, density, average degree, average weight, average clustering coefficient, conductance, and assortativity. These metrics provide a comprehensive characterization of the communities, capturing their structural and functional properties within the network. This approach allows us to evaluate and compare the communities’ characteristics across different modalities systematically.
@@
Figure 7: Cosine similarity between the vectors of the communities detected with Louvain in common with RTW communities. The horizontal line reports the average cosine similarity for each layer’s communities.
@@
For each pair of common labeled communities (as identified in Figure 6a), we calculate the cosine similarity between their metrics. Higher cosine similarity values indicate greater similarity between the communities in terms of their structural and functional properties. Additionally, we compute the mean cosine similarity for the common communities within each layer. The results, presented in Figure 7, reveal that MEN and HST layers exhibit very high levels of cosine similarity, especially in UK, suggesting strong alignment in the structural and functional characteristics of the detected communities. This aligns with the finding that RTW shares a substantial number of communities with both MEN and HST, reinforcing the idea that these modalities capture closely related coordination patterns. In contrast, the URL layer not only has very few communities in common with RTW, but the only shared community also exhibits significant structural differences. This suggests that URL captures a distinct form of coordination, diverging more from retweet-based interactions compared to mention and hashtag. Similar results are obtained also with Infomap, but URL layer exhibits higher values.
@@
A more detailed representation of the common communities’ similarity is presented by Figure 8, which visualizes for UK dataset with Louvain approach, as a starplot, the normalized scores between 0 and 1 of the aforementioned metrics for three examples of couples of common communities. The starplots are color-coded according to the modality to which each community belongs, with a distinct color representing each MONO approach. This visualization allows us to compare the metric scores of matched communities and assess the degree of similarity between them. If two modalities convey similar information, the matched communities will exhibit comparable metric scores, resulting in similar starplot shapes. Conversely, differences in the shapes of the radar plots indicate variations in the information captured by the modalities. This representation effectively highlights the similarities and discrepancies across modes in a clear and intuitive manner. As Figure 6a shows, RPL and RTW do not have communities in common, and for this reason, there is no starplot for comparison. MEN common communities exhibit similar values across all metrics, whereas HST communities are characterized by a high average weight, and their starplot shape slightly differs from that of RTW. Lastly, URL, which shares only one common community with RTW, presents a completely distinct starplot shape, with notably high values for both average weight and assortativity. Similar visualizations are obtained for UK with Infomap approach, and with Louvain and Infomap for IORussia. They are all shown in Appendix A.
@@
Figure 8: Example of metrics comparison of three couple of common communities between RTW and MEN, HST, and URL in UK dataset.
@@
Figure 9: Common communities between RTW and other co-actions in a two-dimensional space using PCA, for UK with Louvain approach. Each community is represented by a vector of metrics, i.e., size, density, average degree, average weight, average clustering coefficient, conductance, and assortativity. These high dimensional vectors are projected into a two-dimensional space using PCA. Communities are represented with distinct point shapes based on the co-action and are linked to their corresponding RTW community. Each RTW community may be associated with multiple communities from different co-actions, with all connected communities sharing the same color for clarity.
@@
The starplot visualization allows to have a quick understanding of the differences between the communities, but it is very detailed and does not provide a global view of the differences. Hence, we represent each community as a feature vector of the aforementioned metrics. Then the feature vectors are projected into a two-dimensional space using Principal Component Analysis (PCA). Figure 9 shows, for UK with Louvain approach, the scatterplot in two PCA dimensions, where each community is linked with the one in common with RTW and identified with a different point shape according to the co-action. Each RTW community may be associated with multiple communities from different co-actions, and the cluster of connected communities are depicted with the same color. The PCA visualization highlights how each cluster of common communities – same color – occupies a distinct region of the plot. This spatial separation suggests that common communities of different INDE are structurally similar to each other while remaining distinct from non-common communities. This reinforces the idea that coordination patterns within common communities are preserved across different modalities. However, an exception emerges with the URL layer, which appears more dispersed and does not conform to the same clustering pattern. This aligns with previous findings that URL is the most distinct and unconventional layer, capturing a different form of coordination compared to the others. Appendix A shows the PCA plots for UK with Infomap approach, and for IORussia with Louvain and Infomap. In these cases, there are some groups of common communities that do not cluster together, indicating that the structural properties of these communities differ more significantly across modalities. This further emphasizes the variability in how different modalities capture coordinated behavior.